# Testing the Fifth Force on Lepton Spins through Neutrino Oscillations

### v1

@@ Abstract
We investigate a fifth force mediated by a light vector boson that couples to lepton spins, characterized by axial-vector couplings to leptons and vector couplings to nucleons. This interaction generates a potential proportional to the inner product of the lepton spin vector and the nucleon-lepton relative velocity vector, a feature extensively explored with precision spin sensors. Employing weak symmetry, we show that left-handed charged lepton couplings naturally extend to left-handed neutrinos, enabling this fifth force to influence neutrino oscillations. For electron-nucleon couplings, we find that solar and reactor neutrino experiments provide comparable constraints to those from spin sensors and surpass them in the short-range fifth force region. For muon-nucleon couplings, neutrino oscillation experiments exclude the fifth force as a viable explanation for the muon $g-2$ anomaly in the context of a vector mediator, tightening the bounds by two orders of magnitude in coupling strength by solar and atmospheric neutrino data. Our results highlight the critical role of neutrino oscillations in probing fifth forces acting across all three generations of lepton spins.
@@ I Introduction
In this work, we explore a fifth force mediated by a light vector boson that couples to lepton spins, extending beyond the scalar mediator studied in our previous work [37]. The mediator is characterized by vector and axial-vector (V-AV) couplings to nucleons and leptons, respectively. We demonstrate that this type of fifth force can be tested through neutrino oscillation experiments. For electron-nucleon couplings, both solar and reactor neutrino experiments impose stringent and complementary constraints compared to spin sensor experiments and surpass them in the short-range fifth force region. For muon-nucleon couplings, our analysis shows that neutrino oscillations rule out the fifth force explanation for the muon $g-2$ anomaly [38, 39] with a vector mediator, tightening the coupling strength constraints by two orders of magnitude. Therefore, neutrino oscillations provide a crucial test for the existence of a fifth force acting on lepton spins.
@@ II Model setup
We consider a $\rm U(1)$ vector gauge boson $A^{\prime}$ couples to leptons ($L$) and nucleons ($N$) through AV and vector couplings, respectively. The general form of its interaction Lagrangian is given by:
@@
$\mathcal{L}_{\rm int}=\bm{g}_{A}^{ij}A^{\prime}_{\mu}\left(-\bar{\nu}_{L}^{i}\gamma^{\mu}\nu_{L}^{j}+\bar{e}^{i}{\gamma}^{\mu}\gamma^{5}e^{j}\right)+g_{V}^{N}A^{\prime}_{\mu}\left(\bar{p}\gamma^{\mu}p+\bar{n}\gamma^{\mu}n\right),$ (2)
where $\bm{g}_{A}=-\bm{g}_{L}=\bm{g}_{R}$ ensures that the coupling to leptons is purely axial-vector. Thus, the vector boson $A^{\prime}$ couples to the charged lepton spins and couples to nucleons with a vector coupling $g_{V}^{N}$.
@@ III Fifth force potential
[…]
@@
Figure 1: The contour plots of ratios between the oscillation probability with and without the fifth force. Upper left panel: the $\nu_{\mu}$ survival probability ratio as a function of neutrino energy $E_{\nu}$ and baseline length $L$, with the black triangle marking the typical neutrino energy and baseline for the T2K experiment. Upper right panel: the $\nu_{\mu}$ survival probability ratio as a function of $E_{\nu}$ and $\cos\theta$, relevant for IceCube experiments. Lower panel: the solar neutrino disappearance probability ratio as a function of $E_{\nu}$ and the coupling $g_{A}g_{V}^{N}$.
@@ IV Neutrino oscillations
For the reactor and accelerator experiments, we assume a constant interaction Hamiltonian $\Delta\mathcal{H}$ during propagation, taking Earth matter density as 2.7 ${\rm g}/{\rm cm}^{3}$ for MSW effects. For ATM neutrinos, we incorporate the variable Earth matter density along the neutrino’s path, segmenting the trajectory to calculate oscillation probability using PREM density data [45]. The production height of atmospheric neutrinos is fixed at 10 km above the Earth surface. For solar neutrinos, we calculate the oscillation probability following [46, 47], with the solar nucleon density from GS98 data [48].
@@
In both the left and central panels of Fig. 1, the ratios display asymptotic behavior at high neutrino energies. In this limit, the vacuum oscillation term becomes small compared to the MSW potential and the fifth force potential. For large values of the product $g_{A}^{\mu\mu}g_{V}^{N}$, the full Hamiltonian contains large contributions in the $11$ and $22$ terms, which suppresses $\nu_{\mu}\to\nu_{e}$ and $\nu_{\mu}\to\nu_{\tau}$ oscillations, resulting in a nearly unchanged survival probability $P^{\rm 5th}_{\mu\mu}$. In contrast, without the fifth force, $\nu_{\mu}\to\nu_{\tau}$ oscillations remain significant, consistent with the SM. Thus, the asymptotic lines in the left and central panels correspond to the minimum probability for two-flavor $\nu_{\mu}\to\nu_{\tau}$ oscillations in the SM scenario. For solar neutrinos in the right panel of Fig. 1, the fifth force effect depends on both the coupling strength and neutrino energy. Stronger couplings and higher energies lead to more significant deviations from SM predictions. For reactor LBL and MBL experiments, since the limits are weaker than other neutrino experiments, we provide their probability ratio plots in the Appendix A.
@@ V Analysis
[…]
@@
where $P_{ee}(E_{i})$ is the theoretical survival probability of electron neutrinos at energy $E_{i}$, and $P_{ee}^{\rm obs}(E_{i})$ and $\delta P_{ee}^{\rm obs}(E_{i})$ are the observed survival probability and its uncertainty. Observed values from BOREXINO  [59] include: $P_{ee}^{\rm obs}$ for $pp$ neutrinos at 0.267 MeV is 0.57 $\pm$ 0.09, for ${}^{7}\text{Be}$ neutrinos at 0.862 MeV is 0.53 $\pm$ 0.05, for $pep$ neutrinos at 1.44 MeV is 0.43 $\pm$ 0.11. For high-energy region ${}^{8}\text{B}$ neutrinos, data points include: at 8.1 MeV, 0.37 $\pm$ 0.08; at 7.4 MeV, 0.39 $\pm$ 0.09; and at 9.7 MeV, 0.35 $\pm$ 0.09. The SNO+SK  [60] analysis provides a survival probability at $10\mathrm{MeV}\left({}^{8}\text{B}\text{~{}neutrinos}\right)$ of $0.308\pm 0.015$, with very small uncertainty. These data, particularly the SNO+SK results, dominate the sensitivity to $g_{A}^{ee}g_{V}^{N}$ and $g_{A}^{\mu\mu}g_{V}^{N}$ due to their relevance at high energies, shown in Fig. 1.
For reactor experiments, we use data from KamLAND [61, 62, 63, 64] and Daya Bay [65, 66, 67, 68]. Since the analysis are similar and their constraints are weaker than the solar, ATM and accelerator neutrino results, we postpone the details of analysis into the Appendix B.
VI Muon g-2
@@
Figure 2: The 95% C.L. constraints on the fifth force model with electron (upper left), muon (upper right) and tau (lower) couplings. The lines show constraints from reactor experiments (Daya Bay [65, 66, 67, 68], KamLAND [61, 62]), solar neutrino experiments (BOREXINO [59], SNO [72], SK [73, 74]), the accelerator experiment T2K [56, 57], and atmospheric neutrino data from IceCube DeepCore [49, 51, 52]. Solid (dashed) lines represent positive (negative) values of the coupling constant $g_{A}g_{V}^{N}$. The existing constraints from spin-sensor experiments are shown in gray, labeled as Heckel 2008 [16] and Clayburn 2023 [19]. The orange hatched region (shaded region) shows parameter space excluded (preferred) by current muon $g-2$ measurements [75, 76, 77, 78].
@@ VIII Results
For $g_{A}^{ee}g_{V}^{N}$ couplings, the primary constraint arises from solar neutrino experiments, which limit $g_{A}^{ee}g_{V}^{N}$ to $[-16.5,1.3]\times 10^{-53}$ in the massless limit and $[-41.6,3.0]\times 10^{-50}\times m_{A^{\prime}}^{2}/(10^{-13}\,\mathrm{eV})^{2}$ in the high-mass region. For positive $g_{A}^{ee}g_{V}^{N}$, solar neutrino limits are stronger than those from Clayburn 2023 study [19]. In the high-mass range, neutrino oscillation constraints scale as $m_{A^{\prime}}^{-2}$ (see Eq. (5)), while spin-sensor constraints scale as $m_{A^{\prime}}^{-3}$, making neutrino oscillations more sensitive for $\lambda\lesssim 10^{5}\,\mathrm{m}$. The reactor experiments KamLAND and Daya Bay provide constraints of $[-13.9,4.3]\times 10^{-50}$ and $[-1.3,2.3]\times 10^{-48}$ in the massless limit.
@@
For $g_{A}^{\mu\mu}g_{V}^{N}$ couplings, the strongest constraints come from ATM neutrino experiments using IceCube data, due to the high neutrino energy. These constraints are $|g_{A}^{\mu\mu}g_{V}^{N}|<1.2\times 10^{-51}\times m_{A^{\prime}}^{2}/(10^{-13}\,\mathrm{eV})^{2}$ at high masses and $|g_{A}^{\mu\mu}g_{V}^{N}|<2.2\times 10^{-52}$ in the massless limit. Solar neutrino and accelerator (T2K) experiments also impose strong constraints, with notable sign dependencies for $g_{A}^{\mu\mu}g_{V}^{N}$, that are not observed in the IceCube results. Typically, if a positive $g^{ee}_{A}g_{V}^{N}$ provides a stronger constraint than the negative, the reverse is true for $g^{\mu\mu}_{A}g_{V}^{N}$ in the same experiment. Our analysis shows that solar, ATM, and accelerator neutrino experiments exclude the fifth force parameter space for $g_{A}^{\mu\mu}g_{V}^{N}$ relevant to the muon $g-2$ anomaly. For $g_{A}^{\tau\tau}g_{V}^{N}$ couplings, the results are similar to $g_{A}^{\mu\mu}g_{V}^{N}$ couplings.
@@ IX Discussions
We examined a fifth force model with lepton spin couplings mediated by a vector boson, using neutrino oscillation experiments as a probe. Our results reveal that neutrino experiments provide constraints on electron-nucleon couplings comparable to those from precision spin sensors, outperforming them in the high-mass region. For muon-nucleon couplings, neutrino oscillations yield constraints two orders of magnitude stronger than those from muon $g-2$ measurements. Tau-nucleon couplings are constrained by neutrino oscillations at a similar level to muon couplings. Our work highlights the capability of neutrino experiments to probe fifth force interactions across all three lepton generations.
@@ Acknowledgement
The work of J.L. is supported by the National Science Foundation of China under Grants No. 12075005, No. 12235001 and No. 12475103. The work of X.P.W. is supported by the National Science Foundation of China under Grants No. 12375095 and the Fundamental Research Funds for the Central Universities.
Appendix A The Oscillation Probability for Reactor neutrino experiments
@@
To analyze the impact of the fifth force on reactor LBL experiments, we fix $m_{A^{\prime}}=10^{-14}\,\text{eV}$ and set $g^{ee}_{A}g^{N}_{V}$ ($g^{\mu\mu}_{A}g^{N}_{V}$) to $-10^{-48}$, respectively. The corresponding probability ratio plots are shown in Fig. 3. In both cases, we observe that for high-energy neutrinos, the ratio reaches maximum deviations when
@@
For the $g^{ee}_{A^{\prime}}\neq 0$ case, the plot exhibits asymmetry around $12V_{m}^{ee}E_{\nu}\cos^{2}\theta_{13}/(2\cos 2\theta_{12}\Delta m_{12}^{2})=$ 1, represented by the horizontal black dot-dashed line. As the energy decreases, the effective mixing angle and mass-squared difference in the fifth force model initially approach the corresponding values in the SM. Specifically, when [62]
@@
Appendix B The analysis for Reactor neutrino experiments
@@
B.1 Reactor Long-baselineexperiments

### v2

@@ Abstract
We investigate a fifth force mediated by a light vector boson that couples to lepton spins, characterized by axial-vector couplings to leptons and vector couplings to nucleons. This interaction generates a potential proportional to the inner product of the lepton spin vector and the nucleon-lepton relative velocity vector, a feature extensively explored with precision spin sensors. Employing weak symmetry, we show that left-handed charged lepton couplings naturally extend to left-handed neutrinos, enabling this fifth force to influence neutrino oscillations. For electron-nucleon couplings, we find that solar and reactor neutrino experiments provide comparable constraints to those from spin sensors and surpass them in the short-range fifth force region. For muon-nucleon couplings, neutrino oscillation experiments exclude the fifth force as a viable explanation for the muon g2g-2 anomaly in the context of a vector mediator, tightening the bounds by two orders of magnitude in coupling strength by solar and atmospheric neutrino data. Our results highlight the critical role of neutrino oscillations in probing fifth forces acting across all three generations of lepton spins.
@@ I Introduction
In this work, we explore a fifth force mediated by a light vector boson that couples to lepton spins, extending beyond the scalar mediator studied in our previous work [37]. The mediator is characterized by vector and axial-vector (V-AV) couplings to nucleons and leptons, respectively. We demonstrate that this type of fifth force can be tested through neutrino oscillation experiments. For electron-nucleon couplings, both solar and reactor neutrino experiments impose stringent and complementary constraints compared to spin sensor experiments and surpass them in the short-range fifth force region. For muon-nucleon couplings, our analysis shows that neutrino oscillations rule out the fifth force explanation for the muon g2g-2 anomaly [38, 39] with a vector mediator, tightening the coupling strength constraints by two orders of magnitude. Therefore, neutrino oscillations provide a crucial test for the existence of a fifth force acting on lepton spins.
@@ II Model setup
We consider a $\rm U(1)^{\prime}$ vector gauge boson $A^{\prime}$ couples to leptons ($L$) and nucleons ($N$) through AV and vector couplings, respectively. The general form of its interaction Lagrangian is given by:
@@
$\mathcal{L}_{\rm int}$ $=\bm{g}_{A}^{ij}A^{\prime}_{\mu}\left(-\bar{\nu}_{L}^{i}\gamma^{\mu}\nu_{L}^{j}+\bar{e}^{i}{\gamma}^{\mu}\gamma^{5}e^{j}\right)$
$+g_{V}^{N}A^{\prime}_{\mu}\left(\bar{p}\gamma^{\mu}p+\bar{n}\gamma^{\mu}n\right),$ (2)
where $\bm{g}_{A}=-\bm{g}_{L}=\bm{g}_{R}$ ensures that the coupling to leptons is purely axial-vector. We assume a flavor-diagonal axial-vector coupling $\bm{g}_{A}^{ij}=a_{i}\delta^{ij}$. To ensure anomaly cancellation, the couplings should satisfy $\sum_{i}a_{i}=\sum_{i}a_{i}^{3}=0$ [40]. This can be achieved via inter-generational cancellation, e.g., $a_{\tau}=-a_{\mu}$, $a_{e}=0$, similar to the $L_{\mu}-L_{\tau}$ model. However, to stay model-independent, we will calculate constraints for one generation at a time.
@@
$\small A^{\prime}_{0}(r)=\frac{-g_{V}^{N}}{2m_{A^{\prime}}r}\int_{0}^{R_{\oplus}}dr^{\prime}n(r^{\prime})r^{\prime}\left(e^{-m_{A^{\prime}}\left|r^{\prime}-r\right|}-e^{-m_{A^{\prime}}\left(r+r^{\prime}\right)}\right),$ (4)
@@
Figure 1: The contour plots of ratios between the oscillation probability with and without the fifth force. Left panel: the $\nu_{\mu}$ survival probability ratio as a function of neutrino energy $E_{\nu}$ and baseline length $L$, with the black triangle marking the typical neutrino energy and baseline for the T2K experiment. Central panel: the $\nu_{\mu}$ survival probability ratio as a function of $E_{\nu}$ and $\cos\theta$, relevant for IceCube experiments. Right panel: the solar neutrino disappearance probability ratio as a function of $E_{\nu}$ and the coupling $g_{A}g_{V}^{N}$.
@@ IV Neutrino oscillations
For the reactor and accelerator experiments, we assume a constant interaction Hamiltonian $\Delta\mathcal{H}$ during propagation, taking Earth matter density as 2.7 ${\rm g}/{\rm cm}^{3}$ for MSW effects. For ATM neutrinos, we incorporate the variable Earth matter density along the neutrino’s path, segmenting the trajectory to calculate oscillation probability using PREM density data [46]. The production height of atmospheric neutrinos is fixed at 10 km above the Earth surface. For solar neutrinos, we calculate the oscillation probability following [47, 48], with the solar nucleon density from GS98 data [49]. Details of our numerical methods can be found in our previous work [37].
@@
In both the left and central panels of Fig. 1, the ratios display asymptotic behavior at high neutrino energies. In this limit, the vacuum oscillation term becomes small compared to the MSW potential and the fifth force potential. For large values of the product $g_{A}^{\mu\mu}g_{V}^{N}$, the full Hamiltonian contains large contributions in the $11$ and $22$ terms, which suppresses $\nu_{\mu}\to\nu_{e}$ and $\nu_{\mu}\to\nu_{\tau}$ oscillations, resulting in a nearly unchanged survival probability $P^{\rm 5th}_{\mu\mu}$. In contrast, without the fifth force, $\nu_{\mu}\to\nu_{\tau}$ oscillations remain significant, consistent with the SM. Thus, the asymptotic lines in the left and central panels correspond to the minimum probability for two-flavor $\nu_{\mu}\to\nu_{\tau}$ oscillations in the SM scenario. For solar neutrinos in the right panel of Fig. 1, the fifth force effect depends on both the coupling strength and neutrino energy. Stronger couplings and higher energies lead to more significant deviations from SM predictions. For reactor LBL and MBL experiments, we provide probability ratio plots in the Appendix.
@@ V Analysis
$\frac{N_{ij}}{N_{ij}^{0}}=\frac{\sum_{a}\iint_{ij}P_{a\rightarrow\mu/\bar{\mu}}\Phi_{a}\,dE\,d\cos\theta}{\sum_{a}\iint_{ij}P^{0}_{a\rightarrow\mu/\bar{\mu}}\Phi_{a}\,dE\,d\cos\theta},$ (8)
@@
where $P_{ee}(E_{i})$ is the theoretical survival probability of electron neutrinos at energy $E_{i}$, and $P_{ee}^{\rm obs}(E_{i})$ and $\delta P_{ee}^{\rm obs}(E_{i})$ are the observed survival probability and its uncertainty. Observed values from BOREXINO  [60] include: $P_{ee}^{\rm obs}$ for $pp$ neutrinos at 0.267 MeV is 0.57 $\pm$ 0.09, for $^{7}\text{Be}$ neutrinos at 0.862 MeV is 0.53 $\pm$ 0.05, for peppep neutrinos at 1.44 MeV is 0.43 $\pm$ 0.11. For high-energy region $^{8}\text{B}$ neutrinos, data points include: at 8.1 MeV, 0.37 $\pm$ 0.08; at 7.4 MeV, 0.39 $\pm$ 0.09; and at 9.7 MeV, 0.35 $\pm$ 0.09. The SNO+SK  [61] analysis provides a survival probability at $10\mathrm{MeV}\left({}^{8}\text{B}\text{~neutrinos}\right)$ of $0.308\pm 0.015$, with very small uncertainty. These data, particularly the SNO+SK results, dominate the sensitivity to $g_{A}^{ee}g_{V}^{N}$ and $g_{A}^{\mu\mu}g_{V}^{N}$ due to their relevance at high energies, shown in Fig. 1.
For reactor experiments, we use data from KamLAND [62, 63, 64, 65] and Daya Bay [66, 67, 68, 69]. Since the analysis are similar and their constraints are weaker than the solar, ATM and accelerator neutrino results, we postpone the details of analysis into the Appendix.
VI Muon g2g-2
@@
Figure 2: The 95% C.L. constraints on the fifth force model with electron (left), muon (center) and tau (right) couplings. The lines show constraints from reactor experiments (Daya Bay [66, 67, 68, 69], KamLAND [62, 63]), solar neutrino experiments (BOREXINO [60], SNO [73], SK [74, 75]), the accelerator experiment T2K [57, 58], and atmospheric neutrino data from IceCube DeepCore [50, 52, 53]. Solid (dashed) lines represent positive (negative) values of the coupling constant $g_{A}g_{V}^{N}$. The red dashed curve is omitted as the IceCube constraints for positive and negative couplings are nearly identical. The existing constraints from spin-sensor experiments are shown in gray, labeled as Heckel 2008 [16] and Clayburn 2023 [19]. The orange hatched region (shaded region) shows parameter space excluded (preferred) by current muon g2g-2 measurements [76, 77, 78, 79].
@@ VIII Results
For $g_{A}^{ee}g_{V}^{N}$ couplings, the primary constraint arises from solar neutrino experiments, which limit $g_{A}^{ee}g_{V}^{N}$ to $16.5,1.3]\times 10^{-53}$ in the massless limit and $41.6,3.0]\times 10^{-50}\times m_{A^{\prime}}^{2}/(10^{-13}\,\mathrm{eV})^{2}$ in the high-mass region. For positive $g_{A}^{ee}g_{V}^{N}$, solar neutrino limits are stronger than those from Clayburn 2023 study [19]. In the high-mass range, neutrino oscillation constraints scale as $m_{A^{\prime}}^{-2}$ (see Eq. (5)), while spin-sensor constraints scale as $m_{A^{\prime}}^{-3}$, making neutrino oscillations more sensitive for $\lambda\lesssim 10^{5}\,\mathrm{m}$. The reactor experiments KamLAND and Daya Bay provide constraints of $13.9,4.3]\times 10^{-50}$ and $1.3,2.3]\times 10^{-48}$ in the massless limit.
@@
For $g_{A}^{\mu\mu}g_{V}^{N}$ couplings, the strongest constraints come from ATM neutrino experiments using IceCube data, due to the high neutrino energy. These constraints are $|g_{A}^{\mu\mu}g_{V}^{N}|<1.2\times 10^{-51}\times m_{A^{\prime}}^{2}/(10^{-13}\,\mathrm{eV})^{2}$ at high masses and $|g_{A}^{\mu\mu}g_{V}^{N}|<2.2\times 10^{-52}$ in the massless limit. Solar neutrino and accelerator (T2K) experiments also impose strong constraints, with notable sign dependencies for $g_{A}^{\mu\mu}g_{V}^{N}$, that are not observed in the IceCube results. Typically, if a positive $g^{ee}_{A}g_{V}^{N}$ provides a stronger constraint than the negative, the reverse is true for $g^{\mu\mu}_{A}g_{V}^{N}$ in the same experiment. Our analysis shows that solar, ATM, and accelerator neutrino experiments exclude the fifth force parameter space for $g_{A}^{\mu\mu}g_{V}^{N}$ relevant to the muon g2g-2 anomaly. For $g_{A}^{\tau\tau}g_{V}^{N}$ couplings, the results are similar to $g_{A}^{\mu\mu}g_{V}^{N}$ couplings.
@@ IX Discussions
We examined a fifth force model with lepton spin couplings mediated by a vector boson, using neutrino oscillation experiments as a probe. Our results reveal that neutrino experiments provide constraints on electron-nucleon couplings comparable to those from precision spin sensors, outperforming them in the high-mass region. For muon-nucleon couplings, neutrino oscillations yield constraints two orders of magnitude stronger than those from muon g2g-2 measurements. Tau-nucleon couplings are constrained by neutrino oscillations at a similar level to muon couplings. Our work highlights the capability of neutrino experiments to probe fifth force interactions across all three lepton generations.
@@ Acknowledgement
The work of J.L. is supported by the National Science Foundation of China under Grants No. 12475103, 12235001 and 12075005. The work of X.P.W. is supported by the National Science Foundation of China under Grants No. 12375095 and the Fundamental Research Funds for the Central Universities.
Appendix A: The Oscillation Probability for Reactor neutrino experiments
@@
To analyze the impact of the fifth force on reactor LBL experiments, we fix $m_{A^{\prime}}=10^{-14}\,\text{eV}$ and set $g^{ee}_{A}g^{N}_{V}$ ($g^{\mu\mu}_{A}g^{N}_{V}$) to $10^{-48}$, respectively. The corresponding probability ratio plots are shown in Fig. 3. In both cases, we observe that for high-energy neutrinos, the ratio reaches maximum deviations when
@@
For the $g^{ee}_{A^{\prime}}\neq 0$ case, the plot exhibits asymmetry around $2V_{m}^{ee}E_{\nu}\cos^{2}\theta_{13}/(2\cos 2\theta_{12}\Delta m_{12}^{2})=1$, represented by the horizontal black dot-dashed line. As the energy decreases, the effective mixing angle and mass-squared difference in the fifth force model initially approach the corresponding values in the SM. Specifically, when [63]
@@
Appendix B: The analysis for Reactor neutrino experiments
@@
IX.1 Reactor Long-baseline experiments